OPTICAL STARK SPECTROSCOPY OF THE Ã^{2}Π- X̃^{2}Σ^{+} BAND OF BaOH

\underline{SARAH E. FREY} AND TIMOTHY C. STEIMLE, Department of Chemistry and Biochemistry, Arizona State University, Tempe,AZ 85287, USA.

Transitions of the Ã^{2}Π- X̃^{2}Σ^{+} band system of barium monohydroxide, BaOH, were observed and recorded from 11483-11485 cm^{−1} and 12041-12044 cm^{−1}. The features were readily identified using the results of the Doppler-limited measurements^{a}. The laser induced fluorescence (LIF) spectrum was analyzed to give optimized field-free excited state parameters. The parameters for the X̃^{2}Σ^{+} state were constrained to the previously determined values^{b}. The permanent electric dipole moments for the X̃^{2}Σ^{+} and Ã^{2}Π states have been determined from the analysis of the optical Stark spectra for the R_{21}(0.5), Q_{21}(1.5), and R_{2}(0.5) lines. The obtained values were μ(X̃^{2}Σ^{+})= 1.426(38)D and μ(Ã^{2}Π)= 0.477(7) D. The results are compared with predicted values from semi-empirical models and those for CaOH and SrOH^{c}.

aJ. G. Wang, J. D. Tandy and P. F. Bernath J. Mol. Spectrosc. 252, 31 (2008)

bM. A. Anderson, M. D. Allen, W. L. Barclay, Jr, and L. M. Ziurys Chem. Phys. Lett 205, 415 (1993)

cT. C. Steimle, D. A. Fletcher, K. Y. Jung and C. T. Scurlock J. Chem. Phys. 96, 2556 (1992)